An electron has been accelerated from rest through a potential difference of 700 V .

A)What is the electron's kinetic energy, in electron volts?

B)What is the electron's speed?
Answer #1

According to the question we need to find the electron kinetic energy and the electron speed.

Potential diff = 700 V

Electron kinetic energy in eV = 700 eV ======PART A)

v = sqrt[2*KE/m] = sqrt ( 2 * 1.1214 * 10 ^ -16 / 9.109x10^-31 ) = 1.5691336 * 10 ^ 7 m/s ======PART B)